# Signatures of Rapidly Rotating Stars with Chemically Homogeneous Evolution in the First Galaxies

### v1

@@ Abstract
The James Webb Space Telescope (JWST) has revealed an unexpectedly high abundance of UV luminous galaxies at redshifts $z\gtrsim 10$, challenging ‘standard’ galaxy formation models. This study investigates the role of rapidly rotating (massive) stars undergoing chemically homogeneous evolution (CHE) in reconciling this potential tension. These stars are more compact, hotter, and exhibit enhanced UV emission. We find that the rest-frame UV luminosity of star-forming galaxies can be significantly enhanced by a factor of $\sim 3-6$ when CHE stars above a minimum initial mass of $m_{\star,\min}^{\rm CHE}\sim 2-10\ \rm M_{\odot}$ account for more than half of the total stellar mass following a Salpeter initial mass function. As a result, the UV luminosity functions observed at $z\sim 12-16$ can be reproduced with less extreme values of star formation efficiency and UV luminosity stochastic variability. Our results highlight the potential of CHE in explaining the UV-bright galaxy populations detected by JWST and call for future work to explore the broader astrophysical implications of CHE and its associated phenomena in the early universe, such as gamma-ray bursts, compact object binaries, and metal enrichment.
@@ 1 Introduction
Early results from the JWST have challenged the standard model of galaxy formation in the high-$z$ universe (Boylan-Kolchin, 2023), with the discovery of a surprising abundance of UV bright galaxies at $z\gtrsim 10$ (e.g., Finkelstein et al., 2023, 2024; Labbé et al., 2023; Adams et al., 2024; Donnan et al., 2024). Two broad classes of explanations have been discussed, modifications to the underlying $\Lambda$CDM model of cosmological structure formation (e.g., Liu & Bromm, 2022; Shen et al., 2024), or models that vary assumptions in the standard picture of star formation. The second (baryonic) category explores scenarios that either increase the stellar mass/star formation rate in a given dark matter host halo, by boosting the star formation efficiency (SFE, e.g., Inayoshi et al., 2022; Dekel et al., 2023) or variability/burstiness of UV emission/star formation (e.g., Shen et al., 2023), or those that enhance the stellar light-to-mass ratio [^1] Alternatively, strong radiation-pressure driven outflows could remove UV dust attenuation during the early evolution of high-$z$ galaxies (e.g., Ferrara, 2024).. An example for the latter explores a top-heavy initial mass function (IMF; e.g., Cueto et al., 2024; Trinca et al., 2024; Ventura et al., 2024), as predicted for the metal-poor environments in the early universe (e.g., Liu et al., 2024).
@@
Thus invoking a top-heavy IMF or an increased formation efficiency may, however, only boost star formation locally, but suppress it on galactic scales, by triggering disruptive stellar feedback, such as strong supernova (SN) activity (e.g., Jeong et al., 2024). Here we propose a different (baryonic) solution for the early UV-bright galaxy problem: The stellar light-to-mass ratio is boosted because of a population of stars that undergo chemically homogeneous stellar evolution (CHE), which become more compact and hotter than normal stars. Such CHE stage could be realized for rapidly rotating stars, where their structure adapts by triggering strong internal mixing currents (e.g., Maeder, 1987; Yoon & Langer, 2005; Cantiello et al., 2007; Yoon et al., 2012; Maeder & Meynet, 2012; Song et al., 2016; Mandel & de Mink, 2016; de Mink & Mandel, 2016; Marchant et al., 2017; du Buisson et al., 2020; Riley et al., 2021). Rotation speeds close to the break-up value have been predicted for the first Population III (Pop III) stars, formed in the unique metal-free conditions [^2] Magnetic fields arising from dynamo activity in the primordial protostellar disks could complicate the situation, leading to a possible bifurcation into cases of rapid and slow rotation (Hirano & Bromm, 2018). of dark matter minihalos (Stacy et al., 2011, 2013). Similarly, metal-poor Population II (Pop II) stars may be able to retain their natal rotation speeds, because of their significantly reduced mass loss due to weakened radiatively-driven stellar winds (e.g., Vink et al., 2001). Recent direct observations of select O-type stars in nearby metal-poor dwarf galaxies suggest that stellar winds may be even weaker than what was previously predicted by theory (Gormaz-Matamala et al., 2024; Telford et al., 2024).
@@
The impact of a CHE-boosted production of H, He, and possibly He+, ionizing radiation from Pop III stars on the early stages of reionization was explored in a previous study (Sibony et al., 2022), placing constraints on the fraction of Pop III stars that undergo CHE (see also Liu et al. in prep.). The reionization modeling, however, is subject to the uncertain escape fraction of ionizing radiation from the galaxy host halos, whereas any constraints from the non-ionizing UV continuum are more direct indications of the underlying stellar populations. We note that there have been recent indications for the metal enrichment from rapidly rotating stars in the first galaxies, inferred from their chemical abundance patterns with deep JWST spectroscopy. Prominent among them are high nitrogen-to-oxgen abundance ratios, e.g., in GN-z11 and RXCJ2248-ID (Bunker et al., 2023; Topping et al., 2024), which may be explained with models of rapidly rotating, massive stars (e.g., Tsiatsiou et al., 2024; Nandal et al., 2024). Another nucleosynthetic signature of rapidly rotating stars is the proposed boost in carbon production (Liu et al., 2021; Jeena et al., 2023), providing a possible explanation for strong carbon emission lines detected in $z\gtrsim 12$ galaxies (e.g., D’Eugenio et al., 2024) and carbon-enhanced extremely metal-poor stars in the Milky Way (e.g., Zepeda et al., 2023). We will therefore explore in the following whether a rapidly rotating subset of stars under CHE within the first galaxies can account for the emerging JWST phenomenology of UV bright galaxies in the first few hundred million years of cosmic history.
@@ 2.1 Stellar Evolution
The stellar evolution models for stars evolving homogeneously are based on $n=3$ polytropes with a time-varying mass fraction of hydrogen, as detailed in Sibony et al. (2022). In that work Pop III stars with initial masses ($m_{\star}$) between 9 and 300 M were computed. Here we complement our previous grid by adding tracks for lower initial mass stars with masses between 1.7 and 7 M. As a comparison sample, we use the stellar tracks for non-rotating Pop III stars in the same mass range ($1.7-300\ \rm M_{\odot}$) from Murphy et al. (2021). Here, we only consider metal-free stellar evolution models for simplicity, while freely varying the IMF parameters (see Sec. 2.2 below). The difference in rest-frame UV ($\lambda\sim 1500\ \rm\AA$) luminosity when changing the metallicity (from $Z=0$ to 0.02) under a fixed Kroupa (2001) IMF is minor (within a factor of $\sim 2$) for a typical star formation timescale $\sim 100$ Myr, according to the Yggdrasil stellar population synthesis models (Zackrisson et al., 2011).
Homogeneous evolution requires that the timescale for chemical mixing is much smaller than the nuclear timescale during the Main-Sequence (MS) phase. The mixing timescale scales as $R^{2}/D_{\rm mix}$, where $R$ is the stellar radius, and $D_{\rm mix}$ the diffusion coefficient describing the transport of the chemical elements. Maeder & Meynet (2000, see their paragraph 3.1) showed that the ratio of the mixing timescale to MS lifetime approximately varies as $1/m_{\star}$ for stars more massive than 15 M. This indicates that when the mass decreases, the mixing timescale increases faster than the MS lifetime. Thus, we can expect to reach a lower initial mass limit, $m_{\star,\min}^{\rm CHE}$, below which homogeneous evolution will not be obtained. The precise limit depends on the initial rotation and metallicity, with lower values when the initial rotation increases and when the metallicity decreases (this is due at least in part to the fact that at lower metallicities, stars are more compact).
@@
Figure 1: Rest-frame UV light curve: evolution of UV luminosity normalized by the total stellar mass for a single-age stellar population. The left (right) panel shows the results for a Salpeter (top-heavy) IMF, $dN/dm_{\star}\propto m_{\star}^{-\alpha}$, with a slope $\alpha=2.35$ (1) and a mass range $m_{\star}\in[1.7,300]\ \rm M_{\odot}$ for both normal and CHE stars. The mass fraction of CHE stars $f_{\rm CHE}$ is varied between 0 and 1, with lighter colors denoting larger $f_{\rm CHE}$.
@@
This limit also depends on the physics used for chemical mixing and for the transport of angular momentum. These two transport mechanisms are intimately related since the gradients of chemical composition impact angular momentum transfer, which in turn influences the transport of the chemical elements. For this work, since we do not use a peculiar physical mechanism to obtain the homogeneous tracks, we simply consider the value of this limit as a free parameter. Most of the stellar models computed so far have focused on the massive stellar range down to lower masses between 5 and 10 M (see e.g., Riley et al., 2021), so we may wonder if it is justified here to extend this limit further. As indicated above, the limit depends on the physics of mixing which is still not yet fully understood. On the other hand, fast rotation is likely not the only process that may lead to homogeneous stars. For instance, grazing collisions have been found to result in nearly completely mixed stars producing blue stragglers, as observed in globular clusters (e.g., Benz & Hills, 1987).
2.2 UV Emission
@@
As in Sibony et al. (2022), we consider stars to radiate as black bodies and compute their average spectral luminosity $L_{\star,\nu}(t)$ in a $100$ window around the typical rest-frame UV wavelength of $1500$ along their evolutionary tracks [^3] Using detailed stellar atmosphere models computed by the code tlusty (Hubeny, 1988; Hubeny et al., 2021), we have verified that the relative errors in $L_{\star,\nu}(t)$ introduced by the back-body assumption are within $30$% for most CHE and non-rotating stars with $m_{\star}\sim 3-300\rm\ M_{\odot}$, surface gravities $g\sim 10^{4.1-5.1}\ \rm cm\ s^{-2}$, and effective temperatures $T_{\rm eff}\sim 10^{4.3}-10^{5}\ \rm K$, that dominate the overall rest-frame UV emission. . We then consider a mixed stellar population containing both CHE and non-rotating stars as a single-age starburst. This population is defined by two parameters, the mass fraction $f_{\mathrm{CHE}}\in[0,1]$ of CHE stars, and the minimal stellar mass $m_{\star,\min}^{\rm CHE}\in[1.7,20\,M_{\odot}]$ for CHE to occur. For a given choice of $(f_{\mathrm{CHE}},m_{\star,\min}^{\rm CHE})$, we first generate two stellar populations following a power-law IMF, $dN/dm_{\star}\propto m_{\star}^{-\alpha}$, with a slope of $\alpha=2.35$ (Salpeter) or $\alpha=1$ (top-heavy). The population of CHE (non-rotating) stars is generated between $m_{\star,\min}^{\rm CHE}$ ($m_{\star,\min}=1.7\ \rm M_{\odot}$) and $m_{\star,\max}=300\rm\,M_{\odot}$. For each population, we compute the total luminosity $L_{\nu}(t)$ by summing up the $L_{\star,\nu}(t)$ for all stars that are still alive at time $t$, up to $t=1\,$Gyr, which is then normalized by the total initial stellar mass $M_{\star}$ as $\mathcal{L}_{\nu}(t)\equiv L_{\nu}(t)/M_{\star}$. Finally, we combine the CHE and non-rotating populations so that the mass fraction of CHE stars is $f_{\mathrm{CHE}}$. For illustration, Figure 1 shows the resulting rest-frame UV light curves for $m_{\star,\min}^{\rm CHE}=1.7\,M_{\odot}$ and varying $f_{\mathrm{CHE}}$.

### v2

@@ Abstract
The James Webb Space Telescope (JWST) has revealed an unexpectedly high abundance of UV luminous galaxies at redshifts $z\gtrsim 10$, challenging ‘standard’ galaxy formation models. This study investigates the role of rapidly rotating (massive) stars undergoing chemically homogeneous evolution (CHE) in reconciling this potential tension. These stars are more compact, hotter, and exhibit enhanced UV emission. We find that the rest-frame UV luminosity of star-forming galaxies can be significantly enhanced by a factor of $\sim 3-6$ when CHE stars above a minimum initial mass of $m_{\star,\min}^{\rm CHE}\sim 2-10~{}\rm M_{\odot}$ account for more than half of the total stellar mass following a Salpeter initial mass function. As a result, the UV luminosity functions observed at $z\sim 12-16$ can be reproduced with less extreme values of star formation efficiency and UV luminosity stochastic variability. Our results highlight the potential of CHE in explaining the UV-bright galaxy populations detected by JWST and call for future work to explore the broader astrophysical implications of CHE and its associated phenomena in the early universe, such as gamma-ray bursts, compact object binaries, and metal enrichment.
@@ 1 Introduction
Early results from the JWST have challenged the standard model of galaxy formation in the high-$z$ universe (Boylan-Kolchin, 2023), with the discovery of a surprising abundance of UV bright galaxies at $z\gtrsim 10$ (e.g., Donnan et al., 2023, 2024; Finkelstein et al., 2023, 2024; Harikane et al., 2023; Labbé et al., 2023; Adams et al., 2024). Two broad classes of explanations have been discussed, modifications to the underlying $\Lambda$CDM model of cosmological structure formation (e.g., Liu & Bromm, 2022; Shen et al., 2024), or models that vary assumptions in the standard picture of star formation. The second (baryonic) category explores scenarios that either increase the stellar mass/star formation rate in a given dark matter host halo, by boosting the star formation efficiency (SFE, e.g., Inayoshi et al., 2022; Dekel et al., 2023) or variability/burstiness of UV emission/star formation (e.g., Shen et al., 2023), or those that enhance the stellar light-to-mass ratio (e.g., Donnan et al., 2025). An example for the latter explores a top-heavy initial mass function (IMF; e.g., Inayoshi et al., 2022; Cueto et al., 2024; Trinca et al., 2024; Ventura et al., 2024), as predicted for the metal-poor environments in the early universe (e.g., Liu et al., 2024). UV emission can also be enhanced when strong radiation-pressure driven outflows remove UV dust attenuation during the early evolution of high-$z$ galaxies (e.g., Ferrara, 2024).
@@
Invoking a top-heavy IMF or an increased formation efficiency may, however, only boost star formation locally, but suppress it on galactic scales, by triggering disruptive stellar feedback, such as strong supernova (SN) activity (e.g., Jeong et al., 2024). Here we propose a different (baryonic) solution for the early UV-bright galaxy problem: The stellar light-to-mass ratio is boosted because of a population of stars that undergo chemically homogeneous stellar evolution (CHE), which become more compact and hotter than stars without CHE [^1] Similar but weaker effects can also be achieved by initial helium enhancement (Katz et al., 2024).. Such CHE stage could be realized for rapidly rotating stars, where their structure adapts by triggering strong internal mixing currents (e.g., Maeder, 1987; Yoon & Langer, 2005; Cantiello et al., 2007; Yoon et al., 2012; Maeder & Meynet, 2012; Eldridge et al., 2011; Eldridge & Stanway, 2012; Szécsi et al., 2015, 2022; Eldridge & Stanway, 2016; Song et al., 2016; Mandel & de Mink, 2016; de Mink & Mandel, 2016; Marchant et al., 2017; Cui et al., 2018; Aguilera-Dena et al., 2018; Kubátová et al., 2019; du Buisson et al., 2020; Riley et al., 2021; Ghodla et al., 2023; Dorozsmai et al., 2024). Rotation speeds close to the break-up value have been predicted for the first Population III (Pop III) stars, formed in the unique metal-free conditions [^2] Magnetic fields arising from dynamo activity in the primordial protostellar disks could complicate the situation, leading to a possible bifurcation into cases of rapid and slow rotation (Hirano & Bromm, 2018). of dark matter minihalos (Stacy et al., 2011, 2013). Similarly, metal-poor Population II (Pop II) stars may be able to retain their natal rotation speeds, because of their significantly reduced mass loss due to weakened radiatively-driven stellar winds (e.g., Vink et al., 2001). Recent direct observations of select O-type stars in nearby metal-poor dwarf galaxies suggest that stellar winds may be even weaker than what was previously predicted by theory (Sander et al., 2017, 2018; Gormaz-Matamala et al., 2024; Telford et al., 2024).
@@
The impact of a CHE-boost in the ionizing radiation of H, He, and possibly He+ from Pop III stars on the early stages of reionization was explored in a previous study (Sibony et al., 2022), placing constraints on the fraction of Pop III stars that undergo CHE (see also e.g. Ghodla et al., 2023, and Liu et al. in prep.). The reionization modeling, however, is subject to the uncertain escape fraction of ionizing radiation from the galaxy host halos, whereas any constraints from the non-ionizing UV continuum are more direct indications of the underlying stellar populations. We note that there have been recent indications for the metal enrichment from rapidly rotating stars in the first galaxies, inferred from their chemical abundance patterns with deep JWST spectroscopy. Prominent among them are high nitrogen-to-oxgen abundance ratios, e.g., in GN-z11 and RXCJ2248-ID (Bunker et al., 2023; Topping et al., 2024), which may be explained with models of rapidly rotating, massive stars (e.g., Tsiatsiou et al., 2024; Nandal et al., 2024). Another nucleosynthetic signature of rapidly rotating stars is the proposed boost in carbon production (Liu et al., 2021; Jeena et al., 2023), providing a possible explanation for strong carbon emission lines detected in $z\gtrsim 12$ galaxies (e.g., D’Eugenio et al., 2024) and carbon-enhanced extremely metal-poor stars in the Milky Way (e.g., Zepeda et al., 2023). As the extreme outcome of rotational mixing, CHE leads to the maximal enhancement of carbon production (Jeena et al., 2023), although it is unclear whether nitrogen enhancement can be achieved by stars with CHE, which presumably depends on how the CHE phase starts and ends. We will therefore explore in the following whether a rapidly rotating subset of stars under CHE within the first galaxies can account for the emerging JWST phenomenology of UV bright galaxies in the first few hundred million years of cosmic history.
@@ 2.1 Stellar Evolution
Figure 1: Rest-frame UV light curve: evolution of UV luminosity normalized by the total stellar mass for a single-age stellar population. The left (right) panel shows the results for a Salpeter (top-heavy) IMF, $dN/dm_{\star}\propto m_{\star}^{-\alpha}$, with a slope $\alpha=2.35$ (1) and a mass range $m_{\star}\in[1.7,300]~{}\rm M_{\odot}$ for both non-rotating and CHE stars. The mass fraction of CHE stars $f_{\rm CHE}$ is varied between 0 and 1, with lighter colors denoting larger $f_{\rm CHE}$.
The stellar evolution models for stars evolving homogeneously are based on $n=3$ polytropes with a time-varying mass fraction of hydrogen, as detailed in Sibony et al. (2022). In that work Pop III stars with initial masses ($m_{\star}$) between 9 and 300 M were computed. Here we complement our previous grid by adding tracks for lower initial mass stars with masses between 1.7 and 7 M. As a comparison sample, we use the stellar tracks for non-rotating Pop III stars in the same mass range ($1.7-300~{}\rm M_{\odot}$) from Murphy et al. (2021). Here, we only consider initially metal-free stellar evolution models for simplicity, while freely varying the IMF parameters (see Sec. 2.2 below). We focus on the Main-Sequence (MS) phase where most UV photons are produced and ignore mass loss from stellar winds, which is minor for metal-poor stars (e.g., Hainich et al., 2015; Kubátová et al., 2019). For metal-rich massive stars where winds are strong, the UV emission will be overestimated in our models. This renders our results as optimistic estimates. In fact, for non-rotating stars, the difference in rest-frame UV ($\lambda\sim 1500~{}\rm\AA$) luminosity made by changing the initial metallicity from $Z=0$ to $Z=0.02$ under a fixed Kroupa (2001) IMF is small (within a factor of $\sim 2$) for a typical star formation timescale $\sim 100$ Myr, according to the Yggdrasil stellar population synthesis models (Zackrisson et al., 2011). For CHE stars, Kubátová et al. (2019) found that changing the wind parameters (mass loss rate and clumping factor) has negligible effects on the UV spectrum around $\lambda\sim 1000-3000\ \rm\AA$ during MS with helium mass fractions $Y\lesssim 0.98$ for $m_{\star}\sim 20-130\ \rm M_{\odot}$ in the metal-poor regime ($Z=0.0002$), although the impact of winds can be stronger at higher metallicities. However, the impact is expected to be smaller than the case of non-rotating stars since the strong surface enhancement of helium is the dominant factor in CHE, as implied by the smaller effects of metallicity on the properties of pure helium stars (see, e.g., figs. 5 and 6 in Iorio et al., 2023).
Homogeneous evolution requires that the timescale for chemical mixing is much smaller than the nuclear timescale during the MS phase. The mixing timescale scales as $R^{2}/D_{\rm mix}$, where $R$ is the stellar radius, and $D_{\rm mix}$ the diffusion coefficient describing the transport of the chemical elements. Maeder & Meynet (2000, see their paragraph 3.1) showed that the ratio of the mixing timescale to MS lifetime approximately varies as $1/m_{\star}$ for stars more massive than $15~{}\rm M_{\odot}$. This indicates that when the mass decreases, the mixing timescale increases faster than the MS lifetime. Thus, we can expect to reach a lower initial mass limit, $m_{\star,\min}^{\rm CHE}$, below which homogeneous evolution will not be obtained. The precise limit depends on the initial rotation and metallicity, with lower values when the initial rotation increases and when the metallicity decreases (see, e.g., fig 4 in Brott et al., 2011), which is due at least in part to the fact that at lower metallicities, stars are more compact.
@@
This limit also depends on the physics used for chemical mixing and for the transport of angular momentum. These two transport mechanisms are intimately related since the gradients of chemical composition impact angular momentum transfer, which in turn influences the transport of the chemical elements. For this work, since we do not use a peculiar physical mechanism to obtain the homogeneous tracks, we simply consider the value of this limit as a free parameter. Most of the stellar models computed so far have focused on the massive stellar mass range down to lower masses between 4 and 10 M (see, e.g., Yoon et al., 2006; Brott et al., 2011; Szécsi et al., 2015; Riley et al., 2021), so we may wonder if it is justified here to extend this limit further. As indicated above, the limit depends on the physics of mixing which is still not yet fully understood. On the other hand, fast rotation is likely not the only process that may lead to homogeneous stars. For instance, grazing collisions have been found to result in nearly completely mixed stars producing blue stragglers, as observed in globular clusters (e.g., Benz & Hills, 1987).
2.2 UV Emission and stellar populations
@@
As in Sibony et al. (2022), we consider stars to radiate as black bodies and compute their average spectral luminosity $L_{\star,\nu}(t,m_{\star})$ in a $100$ window around the typical rest-frame UV wavelength of $1500$ along their evolutionary tracks. We have verified using detailed stellar atmosphere models computed by the code tlusty (Hubeny, 1988; Hubeny et al., 2021) that the relative errors in $L_{\star,\nu}(t,m_{\star})$ introduced by the back-body assumption are within $30$% for most CHE and non-rotating stars with $m_{\star}\sim 3-300\rm~{}M_{\odot}$, surface gravities $g\sim 10^{4.1-5.1}~{}\rm cm~{}s^{-2}$, and effective temperatures $T_{\rm eff}\sim 10^{4.3}-10^{5}~{}\rm K$, that dominate the overall rest-frame UV emission. We also compare our results with those from Kubátová et al. (2019) for CHE stars with $Z=0.0002$ and $m_{\star}\sim 20-130\ \rm M_{\odot}$, whose spectra are calculated using the atmosphere code powr (e.g., Sander et al., 2015) from the evolutionary tracks computed by Szécsi et al. (2015). The discrepancy is also within 30% during MS ($Y\lesssim 0.98$), mainly resulting from the different evolutionary tracks. When $T_{\rm eff}$ is controlled, the difference caused by the black-body assumption alone is within 7%. It is shown below that our conclusions are robust despite these uncertainties.
We consider a mixed stellar population containing both CHE and non-rotating stars as a single-age starburst. This population is defined by two parameters, the mass fraction $f_{\mathrm{CHE}}\in[0,1]$ of CHE stars, and the minimal stellar mass $m_{\star,\min}^{\rm CHE}\in~{}[1.7,20]\rm~{}M_{\odot}$ for CHE to occur. For a given choice of $(f_{\mathrm{CHE}},m_{\star,\min}^{\rm CHE})$, we first generate two stellar populations following a power-law IMF, $dN/dm_{\star}\propto m_{\star}^{-\alpha}$, with a slope of $\alpha=2.35$ (Salpeter, 1955) or $\alpha=1$ (top-heavy). The population of CHE stars is generated between $m_{\star,\min}^{\rm CHE}$ and $m_{\star,\max}=~{}300\rm~{}M_{\odot}$, while non-rotating stars span a fixed mass range from $m_{\star,\min}=1.7~{}\rm M_{\odot}$ to $m_{\star,\max}=300\ \rm M_{\odot}$. For each population, we compute the total luminosity $L_{\nu}(t)$ by summing up the $L_{\star,\nu}(t,m_{\star})$ for all stars that are still alive at time $t$, up to $t=1~{}$Gyr, which is then normalized by the total initial stellar mass $M_{\star}$ as $\mathcal{L}_{\nu}(t)\equiv L_{\nu}(t)/M_{\star}$. Finally, we combine the CHE and non-rotating populations so that the mass fraction of CHE stars is $f_{\mathrm{CHE}}$. For illustration, Figure 1 shows the resulting rest-frame UV light curves for $m_{\star,\min}^{\rm CHE}=~{}1.7\rm~{}M_{\odot}$ and varying $f_{\mathrm{CHE}}$.
@@
Figure 2: Rest-frame UV luminosity per unit SFR $\eta_{\rm UV}$ for a top-hat star formation history within a timescale $t_{\rm SF}=100~{}\rm Myr$. As in Fig. 1, the left and right panel show the results for Salpeter and top-heavy IMFs, with $\alpha=2.35$ and $1$, respectively. While the mass range of non-rotating stars is fixed as $m_{\star}\in[1.7,300]~{}\rm M_{\odot}$, we plot $\eta_{\rm UV}$ as a function of $m_{\star,\min}^{\rm CHE}$ for $f_{\rm CHE}\in[0,1]$, where $f_{\rm CHE}$ increases for lighter colors (from bottom to top). The dotted curve with errorbars shows the maximum $\eta_{\rm UV}$ in the alternative population model with a fixed overall IMF (corresponding to the maximum CHE mass fraction $f_{\rm CHE}^{\max}$), which is estimated to be in between $\eta_{\rm UV}(f_{\rm CHE}=f_{\rm CHE}^{\max})$ and $f_{\rm CHE}^{\max}\eta_{\rm UV}(f_{\rm CHE}=1)$. We also show the reference value $\eta_{\rm UV}\simeq 8.70~{}\times~{}10^{27}$ from Madau & Dickinson (2014) with the dashed line, which is very close to our result for $f_{\rm CHE}=0$ given $\alpha=2.35$. The shaded regions show the results for Pop III stars with extremely top-heavy IMFs from Inayoshi et al. (2022).